A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1890–1899

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

&c [37, 343].

13th [405, 406]. 14th [364, 363]. 15th [135, 134]. 17th [80, 314, 315, 79]. 1890 [34, 31]. 1891 [79]. 1892 [134]. 1894 [227]. 1896 [315]. 1897 [363]. 1898 [405]. 18th [32, 31].

20th [228, 227]. 21st [184]. 23rd [66]. 24th [21]. 25th [261]. 26th [218, 302]. 27th [170]. 28th [121, 393]. 29th [350].

30th [437]. 31st [66, 121, 170, 218, 261, 302, 350, 393, 437, 21].

5th [295].

9 [66, 121, 170, 218, 261, 302, 350, 393, 437, 21].

Abatement [23]. Abroad [338]. Accidents [367]. According [136]. Accounts [92]. Accumulations [2]. Accuracy [366]. Acts [44]. Additions [11, 20, 29, 41, 53, 65, 77, 89, 108, 120, 132, 145, 156, 169, 181, 193, 206, 217, 225, 239, 251, 260, 270, 280, 290, 301, 312, 323, 335, 349, 360, 370, 382, 392, 402, 413, 426, 436, 446, 459]. Address [405, 229, 134, 183, 316, 363, 82, 227, 124, 34, 314, 31, 79, 186, 273]. Adelphi [66, 121, 170, 218, 261, 302, 350, 393, 437, 21]. Affairs [273]. After [363]. Age [208, 404, 353]. Age-Incidence [353]. Aged [94, 81]. Ages [14, 408]. Agricultural [161, 151, 223, 200, 213, 278, 275, 310, 321, 344, 358, 368, 345, 389, 400, 411, 433, 443, 457, 409, 418, 434, 442, 99, 221, 242, 93, 304]. Agriculture [299, 337, 92, 357]. Alcohol [4]. Alien [191, 241]. Alleged [13]. also [221]. America [211]. among [44, 284]. Annual [405, 22, 66, 122, 121, 171, 170, 219, 218, 261, 302, 350, 394, 393, 437, 351, 438, 262, 21, 67, 303]. Anthropometric [18]. Application [296, 439, 453]. Applications [355, 376, 397]. Approaches [198]. April [295]. Area [114]. Article [205]. Articles [289]. Artisan [45]. Aspect [398, 429]. Aspects [374, 354, 34, 56]. Association [126, 187, 235, 229, 316, 82, 124, 34, 186, 273]. Asymmetrical [222]. Australasia [17]. Australasian [416]. Australia [165]. Austria [118, 140]. Austria-Hungary [118]. Average [253, 366, 356]. Averages [68].

Bankruptcies [268]. Banks [275, 338, 304]. Based [183]. be [277]. Bearing [384, 197, 263]. Belgium [275, 448]. Between [253, 189, 222, 227, 230, 15, 408, 246, 366]. Biddulph [340]. Bills [125]. Bimetallic [374]. Bimetallism [263]. Birmingham [256]. Birth [396, 420, 408]. Block [45]. Board [212]. Booth [94]. Brigade [7, 49, 102, 152, 203, 248, 287, 332, 379, 423]. Britain [223, 278, 246, 310, 321, 358, 368, 400, 411, 443, 110]. Britian [197]. British [126, 187, 235, 34, 186, 70, 84, 247, 299, 372, 229, 316, 82, 124, 210, 456, 264, 273]. Budapest [233]. Budgets [160]. Buildings [45]. Bulgaria [141].

Caird [91]. Calculation [356]. Calculus [355, 376, 397]. can [277]. Canadian [26]. Canal [27]. Cancer [399]. Capital [2]. Cardiff [82]. Case [221, 256]. Caused [140]. Causes [284, 428]. Census [17, 59, 69, 137, 129, 143, 117, 167, 212, 237, 277, 319, 309, 329, 432, 452, 96, 388, 183, 210, 57, 39, 231]. Census-Taking [210]. Censuses [70, 84]. Century [74]. Certain [408]. Ceylon [142]. Chance [25, 35]. Changes [253, 342, 428]. Charges [317, 305]. Charitable [56]. Charles [113, 134, 183]. Chicago [201]. Chief [43, 440]. Chiefly [428]. Child [396, 420, 408]. Child-Birth [396, 420, 408]. Children [408, 149, 284]. China [188, 419]. Christiania [444]. City [256]. Civil [410]. Civilisation [63, 328]. Class [449]. Classes [48]. Classical [186]. Classification [94, 81]. Coal [320, 33, 363, 230]. Coinage [37]. Collection [384]. College [92, 347]. Colleges [242]. Colonial [440]. Colonies [70, 84]. Colored [60]. Commercial [5, 46, 97, 150, 199, 244, 285, 330, 377, 405, 421]. Commission [357, 352]. Committee [329]. Commodities [188, 245, 6, 47, 98, 159, 202, 286, 331, 378, 422]. Communities [294]. Comparability [214]. Comparative [416, 264]. Comparison [221, 456]. Compensation [367]. Competition [34]. Competitive [25, 35]. Conciliation [24, 15]. Condition [149]. Conditions [209, 337, 221, 284]. Congress [75, 444]. Connected [254]. Conquest [37]. Conseil [178]. Consideration [58]. Constituent [366]. Construction [439, 453]. Consumption [215, 236, 110, 4]. Contagious [44]. Continent [73]. Continental [456]. Contours [234]. Contributions [265, 190, 343, 298]. Corn [257, 337]. Correction [111]. Correlation [373, 189, 222, 365]. Cosmopolitanism [82]. Cost [367]. Council [66, 121, 167, 170, 218, 261, 302, 350, 393, 398, 21]. Countries [70, 214, 33, 138, 264, 4, 185]. County [167, 398]. Couples [408]. Course [372]. Courtney [405]. Creation [86]. Credit [304]. Crime [23]. Criminal [325]. Criminals [18]. Criticisms [420]. Cube [356]. Currency [118, 175, 174, 258].

Curves [293]. Customs [62].

Damage [140]. Danson [387]. Data [296]. Death [91, 326]. Death-Rates [326]. Deaths [396, 408, 420]. Debt [179]. Decade [109]. Decades [428]. December [405, 363, 66, 121, 170, 218, 261, 302, 364, 350, 406, 393, 437, 21]. Decennial [308]. Decimal [258]. Defective [165]. Defects [373, 284]. Defence [43]. Degree [233]. Delivered [227, 405, 134, 363, 314, 31, 79]. Demographic [373]. Demography [154, 75]. Department [166, 148]. Depopulation [172]. Depression [320, 92, 242, 93]. Deputation [212]. Derived [234]. Determination [284]. Development [385]. Developments [314]. Different [175, 138]. Discovery [211]. Discussed [388, 362]. Discussion [94]. Diseases [44]. Dissolution [86]. Distribution [266, 147, 317, 296, 307]. Disturbance [165]. Dominions [210]. Dr. [341]. Drawn [233]. Dulness [284]. During [7, 49, 409, 418, 434, 442, 455, 254, 337, 23, 314, 159, 85, 428, 302]. Duties [62]. Dwellings [45].

Each [127]. Earth [87]. Economic [229, 316, 82, 124, 221, 34, 273, 186]. Economical [10, 19, 28, 40, 52, 64, 76, 88, 106, 119, 131, 144, 155, 168, 180, 192, 216, 224, 238, 205, 250, 259, 269, 279, 300, 311, 322, 289, 334, 348, 359, 369, 381, 391, 401, 412, 425, 435, 445, 458]. Economics [82, 227]. Economy [186]. Edinburgh [124]. Education [209, 3]. Effect [188, 127]. Effects [384, 93]. Efficiency [138]. Eight [121]. Eighth [122]. Elected [127]. Elections [283, 398]. Electrical [232]. Element [25, 35]. Eleventh [231]. Emigration [26, 61, 115, 163]. Empire [372]. Empires [440]. Employed [230]. Employing [234]. Employment [267, 384]. Ended [66, 121, 170, 218, 261, 350, 393, 23, 21, 437]. Ending [66, 121, 170, 218, 261, 302, 350, 393, 21, 437]. England [74, 212, 268, 283, 23, 3, 221, 13, 220, 37, 326, 293, 428]. English [9, 50, 105, 153, 204, 249, 288, 333, 380, 424, 33, 57, 352]. Enumeration [81, 111, 94]. Epidemics [352]. Errata [195]. Esq [314]. Esq. [134, 183]. Essay [154]. Essex [337]. Estimate [233]. Estimates [165, 57]. Ethics [55]. Eventual [127]. Evolution [190, 298, 343]. Examination [33]. Examinations [25, 35]. Excess [415]. Exemplified [337]. Expansion [405]. Expenditure [43, 256]. Experiment [405]. Experiments [123]. Exports [342]. Extent [277, 384]. External [158].

F [126, 187, 235, 34]. F. [31, 79]. F.G.S. [375]. F.R.C.S [31, 79]. F.R.S [375]. Fall [188]. Famine [95]. Farm [337, 173]. Farmer [456]. Farrer [450, 227]. Favour [186]. Fellows [127]. Female [130]. Fifteen [127]. Fifth [437, 438]. Fifty [22, 66, 122, 121, 171, 170, 337, 21, 67, 149, 284]. Fifty-Eight [121]. Fifty-Eighth [122]. Fifty-Ninth [171, 170]. Fifty-Seventh [66, 67]. Fifty-Sixth [22, 21]. Finance [256, 255]. Finances [176]. Financial [66, 121, 170, 218, 246, 261, 302, 350, 393, 437, 21]. Fire [7, 49, 102, 152, 203, 248, 287, 332, 379, 423]. Fires [7, 49, 102, 152, 203, 248, 287, 332, 379, 423, 103]. First [261, 183, 262]. Fisheries [116]. Flag [440]. Food [448, 197]. Foreign [70, 264]. Former [317]. Formulæ [441, 407, 417, 431]. Fourth [394, 393]. France [129, 179, 274, 48, 448, 221, 73, 307, 85, 185]. Frederic [341]. French [178]. Frequency [293]. Frequency-Curves [293]. Friendly [254, 255]. Fund [346]. Future [58].

Galton [297]. Gazette [166]. General [22, 66, 122, 121, 171, 170, 219, 218, 277, 261, 302, 350, 394, 393, 437, 351, 438, 262, 21, 67, 303, 308]. Geometrical [454]. German [8, 51]. Germany [275, 319, 448, 367]. Girls [267]. Gladstone [386]. Gold [211, 430, 37, 272]. Gothenburg [164]. Government [212]. Great [223, 278, 246, 310, 321, 358, 368, 400, 411, 443, 197, 110]. Growth [221, 14].

H [405, 363]. Hail [140]. Harvest [247]. Held [66, 121, 170, 218, 350, 229, 316, 124, 21, 273, 261, 302, 393, 437, 82, 34, 186]. Highways [197, 198]. Hippophagy [73]. History [5, 46, 97, 150, 199, 244, 285, 330, 377, 421, 293]. Holdings [345]. Home [173, 338, 44]. Homicide [128]. Hon [405, 386, 363, 227]. Hospital [93]. Hours [138]. Hundred [409, 418, 434, 442, 455, 149]. Hungarian [72]. Hungary [118]. Hygiene [354, 75].

Identification [18]. III [434]. Illustrated [283, 256]. Illustrations [454, 255]. Immigration [26, 61, 115, 191, 163, 241]. Imperial [39]. Imports [342, 415]. Inaugural [134, 363, 227, 314, 31]. Incidence [274, 317, 353]. Inclosure [339]. Income [274, 48]. Incomes [266]. Increased [399]. Increasing [13]. Index [38]. India [59, 327, 390, 96, 147, 209, 318, 39]. Indian [174, 410]. Individual [274]. Industrial [384, 24, 15, 292]. Industry [230]. Infant [196]. Influence [62, 63]. Information [1, 30, 42, 78, 90, 133, 146, 182, 194, 226, 240, 271, 281, 313, 324, 361, 371, 403, 414, 447]. Inland [129]. Inquiry [277, 183, 243, 125, 110, 149]. Insanity [13]. Institute [201, 276, 388, 362]. Institutions [254]. Intercensal [428]. Interest [58]. International [139, 75, 201, 276, 444, 388, 362]. Interpretation [325]. Introduction [44, 293]. Investigation [428]. Investigations [173]. Ipswich [273]. Ireland [137, 246]. Iron [33]. Isogens [234]. Italy [275]. IV [442].

J [31, 79]. James [91, 375]. Jevons [363]. John [341, 340, 387, 314]. June [66, 121, 170, 218, 261, 302, 350, 393, 437, 21].

Keleti [113]. Kingdom

[69, 116, 457, 158, 253, 409, 418, 434, 442, 455, 373, 448, 2, 83, 44, 295, 198]. Korosi [234].

Labour [130, 166, 183, 337, 138, 55, 148]. Labourer [45]. Land [74]. Lands [339]. Large [356]. Largest [57]. Last [409, 418, 434, 442, 455, 159, 428]. Latter [317]. Leading [93]. Leeds [34]. Legitimate [233]. Leonard [405]. Library [11, 20, 29, 41, 53, 65, 77, 89, 108, 120, 132, 145, 156, 169, 181, 193, 206, 217, 225, 239, 251, 260, 270, 280, 290, 301, 312, 323, 335, 349, 360, 370, 382, 392, 402, 413, 426, 436, 446, 459]. Life [183, 439, 453, 196]. Life-Tables [453]. Life-Tables-Their [439]. Limitation [127]. Liquor [164]. List [251, 260, 270, 280, 290, 301, 312, 323, 335, 349, 360, 370, 382, 392, 402, 413, 426, 436, 446, 459]. Literature [9, 8, 50, 51, 105, 153, 204, 249, 288, 333, 380, 424]. Litigation [220]. Liverpool [316]. LL.D. [341, 31, 79]. Local [162, 212, 395, 256, 326]. Lock [100]. Lock-Outs [100]. London [7, 49, 66, 102, 121, 167, 152, 170, 203, 218, 248, 261, 287, 309, 302, 332, 350, 379, 393, 423, 437, 398, 183, 395, 21, 93]. Longevity [410]. Lord [227, 450].

M. [113]. M.A [405, 363, 375]. M.A. [340, 314]. M.D [341, 31, 79]. M.P [405]. M.P. [363]. Machine [232]. Made [277, 233]. Magnitudes [356]. Making [277]. Man [343]. Marriage [14, 408].

Marriage-Ages [14]. Marriage-Rates [14]. Martin [340]. Mathematical [441, 407, 417, 431, 190, 298, 343]. Matters [388]. Measure [328]. Medals [112]. Medical [56]. Meeting [388, 362, 34, 22, 66, 122, 121, 171, 170, 219, 218, 261, 302, 350, 394, 393, 437, 351, 438, 262, 21, 67, 303]. Men [44]. Mental [354, 149, 284]. Method [234, 296, 293]. Methods [314, 284]. Metropolitan [7, 49, 102, 152, 203, 248, 287, 332, 379, 423]. Military [43]. Milk [110]. Mines [83]. Mining [367, 230]. Miscellaneous [355, 376, 397]. Modes [210]. Money [272]. Moral [354]. Morals [227]. Morbidity [136]. Mortality [399, 136, 125]. Mortgages [173]. Mouat [341]. Movement [147, 138, 63, 85]. Movements [339]. Mr. [94, 340, 296, 297]. Municipal [256].

Natality [234, 233]. National [179, 48, 282]. Nationalism [82]. Nations [175]. Nature [68]. Naval [43]. Navigation [129]. Navy [44]. Ninth [171, 170]. Nominal [253]. Norway [164, 266]. Note [58, 75, 39, 419, 297, 347, 298]. Notes [10, 19, 28, 40, 52, 64, 76, 88, 70, 84, 106, 119, 131, 144, 155, 168, 180, 192, 216, 224, 238, 250, 259, 269, 279, 300, 311, 322, 334, 348, 359, 369, 381, 391, 401, 412, 425, 435, 445, 458, 444, 362, 448, 306, 293]. Notification [282]. Nottingham [186]. November [315, 134, 227, 31, 79, 32, 80, 135, 184, 228, 314]. Number [345, 408, 356, 127]. Numbers [38, 230].

Obituary [341, 340]. Observations [317, 233, 284]. Occupation [136]. Occupations [319]. Ocean [197, 198]. Offered [154]. Office [254]. Official [384]. Old [208, 404, 125]. One [337, 149]. Open [339]. Opening [79, 183]. Operation [44]. Ordinary [127]. Other [254, 373, 284, 185, 45]. our [148]. Outs [100]. Oxford [229, 92, 242].

Paper [94, 420]. Papers [58]. Paris [103]. Parks [167]. Parliamentary [283]. Part [409, 418, 455, 407, 428, 434, 442]. Parts [366]. Past [337]. Pauperism [101, 208, 296, 293, 428]. Paupers [94, 81, 111]. Peabody [45]. Pensions [94, 81, 404]. People [183]. Percentiles [296]. Perils [196]. Periodical [12, 54, 107, 157, 207, 252, 291, 336, 383, 427]. Periodicals [205, 289]. Personal [307]. Persons [167]. Petersburg [388, 362]. Phenomena [189, 222]. Physical [354, 149, 284]. Political [186]. Poor [385]. Popular [209, 3]. Population [36, 71, 60, 141, 142, 147, 165, 57, 63, 114, 295, 14, 85, 87]. Portugal [237]. Position [221, 24]. Possessions [264]. Possibility [277]. Powers [43]. Pox [352, 353]. Practical [439, 453, 273]. Preface [87]. Preliminary [70, 84, 75, 59, 69]. Present [295, 37]. Presented [66, 121, 170, 218, 261, 302, 350, 393, 21, 437]. Preservation [339]. President [405, 212, 134, 183, 363, 82, 227, 124, 34, 314, 31, 79]. Prevalence [13]. Price [62]. Prices [188, 257, 245, 327, 390, 342, 243, 337, 430, 6, 47, 98, 159, 202, 286, 331, 378, 422]. Principal [16, 33, 264]. Printers [455]. Prison [55, 139]. Prize [154]. Probabilities [355, 376, 397]. Probable [408, 127]. Proceedings [22, 32, 80, 122, 135, 126, 171, 184, 187, 228, 219, 235, 315, 364, 406, 394, 444, 351, 438, 262, 67, 303]. Produce [161]. Producing [33]. Production [211, 33, 110]. Products [110]. Progress [158, 254, 73, 449]. Property [266, 307]. Proposals [374]. Proposed [329]. Proprietorship [173]. Prospects [209, 24]. Protection [196]. Provision [58].

Quarries [83]. Quarterly [251, 260, 270, 280, 290, 301, 312, 323, 335, 349, 360, 370, 382, 392, 402, 413, 426, 436, 446, 459]. Question [33, 363, 429].

Raiffeisen [275]. Railway [104, 305]. Railways [416, 72]. Rates [317, 83, 14, 305, 326]. Rawson [451, 451]. Reaction [186]. Real [253]. Recorded [284]. Records [337, 165]. Reference [329, 317, 33, 337, 221, 13, 352, 138, 353, 14, 284]. Reform [118, 15]. Reformatory [292]. Regarding [419]. Registrar [308]. Registrar-General [308]. Registration [282]. Registry [254]. Regulating [164]. Relating [449]. Relation [230]. Relations [246, 366, 227, 15, 273]. Relief [385, 56]. Religious [96]. Remarks [258, 297]. Rent [454]. Reorganisation [148]. Repeal [44]. Reply [420]. Report [69, 66, 121, 170, 164, 218, 261, 302, 350, 329, 357, 393, 83, 21, 352]. Representation [441, 283, 407, 417, 431]. Reproductive [298]. Research [346, 314]. Results [183, 57, 149, 234]. Return [59, 83]. Returns [12, 70, 84, 54, 107, 151, 157, 223, 200, 213, 207, 278, 252, 310, 321, 291, 344, 358, 368, 336, 389, 400, 411, 383, 433, 443, 457, 427, 99, 104]. Review [5, 46, 97, 150, 199, 244, 285, 330, 377, 421, 440]. Right [405, 386, 363, 227]. Rooms [66, 121, 170, 218, 261, 302, 350, 393, 437, 21]. Royal [405, 134, 183, 363, 227, 314, 31, 79, 66, 112, 121, 170, 218, 261, 302, 350, 357, 393, 437, 21, 352, 44, 127]. Rural [172]. Russian [95].

Sale [164]. Sanitation [353]. Savings [338]. School [354]. Schools [149, 292]. Science [229, 316, 82, 124, 34, 186, 273]. Scotland [385]. Sea [116]. Second [302, 303]. Section [229, 316, 34, 186, 273, 126, 187, 235, 82, 124]. seen [149, 284]. Selection [298]. September [34]. Service [410]. Session [227, 405, 134, 183, 363, 314, 31, 79]. Sessional [66, 121, 170, 218, 261, 302, 350, 393, 437, 21]. Seven [159]. Seventh [75, 66, 67]. Sex [353]. Sex-Incidence [353]. Showing [408]. Shown [73, 92]. Siam [177]. Sickness [282]. Silver [143, 188, 211, 327, 390, 430, 37, 272]. Since [211, 3, 449, 342, 220]. Sir [91, 451]. Six [149]. Sixth [22, 75, 21]. Sixtieth [219, 218, 34]. Sixty [261, 302, 350, 394, 393, 437, 351, 438, 262, 303]. Sixty-Fifth [437, 438]. Sixty-First [261, 262]. Sixty-Fourth [394, 393]. Sixty-Second [302, 303]. Sixty-Third [350, 351]. Size [345]. Small [352, 353]. Small-Pox [352, 353]. Social [189, 222, 15]. Societies [254]. Society [58, 66, 112, 121, 170, 218, 261, 302, 350, 405, 393, 437, 134, 183, 363, 227, 21, 314, 31, 79, 127, 255]. Some [70, 84, 317, 265, 314, 419, 263, 255, 449, 34]. South [130, 396, 420]. Southern [60]. Special [317, 337, 13, 352, 353, 14, 284]. Specially [221]. Spirituous [164]. Square [356]. St. [388, 362]. Stability [430]. State [94, 81, 123]. States [36, 36, 71, 60, 128, 117, 176, 173, 231]. Station [44]. Statistical [10, 19, 28, 40, 52, 64, 76, 88, 66, 112, 106, 119, 131, 144, 121, 155, 168, 180, 192, 170, 201, 216, 224, 238, 218, 205, 276, 250, 259, 269, 279, 261, 300, 311, 322, 302, 289, 334, 348, 359, 369, 350, 405, 381, 391, 401, 412, 393, 437, 425, 435, 445, 458, 444, 374, 388, 398, 134, 183, 363, 362, 189, 227, 21, 314, 429, 31, 79, 186]. Statistics [140, 215, 268, 267, 441, 229, 214, 208, 409, 418, 434, 442, 455, 373, 384, 316, 82, 43, 265, 306, 294, 407, 417, 431, 227, 236, 124, 109, 23, 416, 13, 352, 385, 220, 34, 353, 73, 325, 139, 45, 264, 419, 273, 110, 263, 449, 347]. Statistique [178]. Storied [398]. Strand [66, 121, 170, 218, 261, 302, 350, 393, 437, 21]. Strikes [100]. Studing [284]. Subjects [362]. Suez [27]. Sugar [429]. Suggestions [258]. Summary [440]. Supérieur [178]. Supplement [308]. Supplementary [306].

Supply [243, 448]. Synchronous [277, 329]. System [164, 72, 282]. Systems [175, 221].

Table [234, 408]. Tables [211, 439, 453, 87]. Tabulating [232]. Taking [210]. Taxation [162, 274, 395, 256]. Taxes [317]. Tea [215, 236]. Teaching [347]. Telegony [343]. Ten [254, 326]. Terminal [305]. Terrace [66, 121, 170, 218, 261, 302, 350, 393, 437, 21]. Territories [36]. Textile [16]. Their [373, 197, 221, 439, 453, 44]. Theory [265, 454, 190, 298, 343, 347, 365]. Third [350, 351]. Thirty [363]. Thousand [149, 284]. Time [37]. Tonnage [109]. Total [127]. Towne [387]. Towns [57]. Trade [177, 320, 318, 372, 214, 158, 440]. Trades [16, 86]. Transfer [317]. Treated [293]. Troops [44]. Trustworthiness [125]. Twenty [23]. Two [428].

Ultimate [44]. Uniform [277]. Uniformity [17]. Unions [86]. United [60, 36, 69, 71, 128, 116, 117, 176, 457, 158, 253, 409, 418, 434, 442, 455, 373, 448, 2, 83, 173, 44, 295, 231, 198]. Universal [329]. University [347]. Unprogressive [294]. Upon [263, 63]. Urban [71]. Urn [398]. Uses [68]. Using [167]. Utilisation [384].

V [455]. Vaccination [352, 353]. Value [58, 188]. Values [74]. Various [214, 4]. Victoria [346]. Vital [45]. Volume [1, 30, 42, 78, 90, 133, 146, 182, 194, 226, 240, 271, 281, 313, 324, 361, 371, 403, 414, 447].

W [386, 451]. Wages [16, 253, 409, 418, 434, 442, 455, 48, 83, 197, 230, 138]. Wales [74, 130, 212, 268, 396, 420, 23, 3, 220, 326, 293]. want [373]. War [221]. Wealth [221, 185]. Wheat [243, 62]. which [277]. White [60]. within [372]. Women [267, 384]. Work [138]. Working [48, 449]. Workmen [160]. Works [10, 19, 28, 40, 52, 64, 76, 88, 106, 119, 131, 144, 155, 168, 180, 192, 216, 224, 238, 250, 259, 269, 279, 300, 311, 322, 334, 348, 359, 369, 381, 391, 401, 412, 425, 435, 445, 458]. World [161, 211, 33, 272].

Year [7, 49, 66, 102, 121, 170, 218, 261, 302, 350, 393, 437, 21, 127, 85]. Years [158, 409, 418, 434, 442, 455, 254, 363, 337, 430, 23, 314, 159, 326]. Yule [296].

Zealand [123]. Zone [72].

References

Anonymous:1890:VIa

[1] Anonymous. Volume information. Journal of the Royal Statistical Society, 53 (1):i–viii, March 1890. ISSN 0952-8385.

                                             URL http://www.jstor.org/stable/
                                             2979520.

Giffen:1890:ACU

[2] Robert Giffen. Accumulations of capital in the United Kingdom in 1875– 85. Journal of the Royal Statistical Society, 53(1):1–49, March 1890. ISSN

                                             0952-8385. URL http://www.jstor.
                                             org/stable/2979521.

Hamilton:1890:PEE

[3] Rowland Hamilton. Popular education in England and Wales since 1882. Journal of the Royal Statistical Society, 53 (1):50–112, March 1890. ISSN 0952-

                                             8385. URL http://www.jstor.org/
                                             stable/2979522.

Yvernes:1890:CAV

[4] E. Yvernes. The consumption of alcohol in various countries. Journal of the Royal Statistical Society, 53(1):113– 127, March 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979523.

Anonymous:1890:CHR

[5] Anonymous. Commercial history and review of 1889. Journal of the Royal Statistical Society, 53(1):127–141, March 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979524.

Sauerbeck:1890:PC

[6] A. Sauerbeck. Prices of commodities in 1888 and 1889. Journal of the Royal Statistical Society, 53(1):141–145, March 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979525.

Anonymous:1890:FLD

[7] Anonymous. Fires in London during the year 1889, and the Metropolitan Fire Brigade. Journal of the Royal Statistical Society, 53(1):145–150, March 1890. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979526.

Anonymous:1890:GL

[8] Anonymous. German literature in 1888 and 1889. Journal of the Royal Statistical Society, 53(1):150, March 1890. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979527.

Anonymous:1890:EL

[9] Anonymous. English literature in 1889. Journal of the Royal Statistical Society, 53(1):151–153, March 1890. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979528.

Anonymous:1890:NESa

[10] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 53(1):153– 163, March 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979529.

Anonymous:1890:ALa

[11] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 53(1):163–176, March 1890. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979530.

Anonymous:1890:PR

[12] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 53(1): 177–200, March 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979531.

Humphrey:1890:SIE

[13] Noel A. Humphrey. Statistics of insanity in England, with special reference to its alleged increasing prevalence. Journal of the Royal Statistical Society, 53(2): 201–252, June 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979408.

Ogle:1890:MRM

[14] William Ogle. On marriage-rates and marriage-ages, with special reference to the growth of population. Journal of the Royal Statistical Society, 53(2):253–289, June 1890. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979409.

Price:1890:RBI

[15] L. L. Price. The relations between industrial conciliation and social reform. Journal of the Royal Statistical Society, 53(2):290–302, June 1890. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979410.

Anonymous:1890:WPT

[16] Anonymous. Wages in the principal textile trades. Journal of the Royal Statistical Society, 53(2):303–309, June 1890. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979411.

Anonymous:1890:UCA

[17] Anonymous. The uniformity of the census of Australasia in 1891. Journal of the Royal Statistical Society, 53(2):309–323, June 1890. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979412.

Anonymous:1890:AIC

[18] Anonymous. Anthropometric identification of criminals. Journal of the Royal Statistical Society, 53(2):323–326, June 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979413.

Anonymous:1890:NESb

[19] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 53(2):326–337, June 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979414.

Anonymous:1890:ALb

[20] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 53(2):337–356, June 1890. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979415.

Hendriks:1890:RCF

[21] Frederick Hendriks. Report of the Council for the financial year ended 31st December, 1889, an for the sessional year ending 24th June, 1890, presented at the fifty-sixth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 24th of June, 1890. Journal of the Royal Statistical Society, 53(3):357–372, September 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979442.

Anonymous:1890:PFS

[22] Anonymous. Proceedings of the fiftysixth Annual General Meeting. Journal of the Royal Statistical Society, 53(3): 373–376, September 1890. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979443.

Grosvenor:1890:SAC

[23] George Grosvenor. Statistics of the abatement in crime in England and Wales, during the twenty years ended 1887–88. Journal of the Royal Statistical Society, 53(3):377–419, September 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979444.

Price:1890:PPI

[24] L. L. Price. The position and prospects of industrial conciliation. Journal of the Royal Statistical Society, 53(3):420– 459, September 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979445.

Edgeworth:1890:ECCa

[25] F. Y. Edgeworth. The element of chance in competitive examinations. Journal of the Royal Statistical Society, 53(3):460– 475, September 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979446.

Anonymous:1890:CIE

[26] Anonymous. Canadian immigration and emigration. Journal of the Royal Statistical Society, 53(3):476–491, September 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979447.

Anonymous:1890:SC

[27] Anonymous. The Suez Canal. Journal of the Royal Statistical Society, 53(3):491– 494, September 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979448.

Anonymous:1890:NESc

[28] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 53(3):494–508, September 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979449.

Anonymous:1890:ALc

[29] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 53(3):508–520, September 1890. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979450.

Anonymous:1890:VIb

[30] Anonymous. Volume information. Journal of the Royal Statistical Society, 53 (4):725–735, December 1890. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979542.

Mouat:1890:IAF

[31] F. J. Mouat. The inaugural address of F. J. Mouat, M.D., F.R.C.S., LL.D., President of the Royal Statistical Society. Session 1890–91. Delivered 18th November, 1890. Journal of the Royal Statistical Society, 53(4):521–556, December 1890. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979543.

Anonymous:1890:PN

[32] Anonymous. Proceedings on the 18th November, 1890. Journal of the Royal Statistical Society, 53(4):557– 560, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979544.

Chisholm:1890:ECI

[33] George G. Chisholm. An examination of the coal and iron production of the principal coal and iron producing countries of the world, with reference to the English coal question. Journal of the Royal Statistical Society, 53(4):561– 611, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979545.

Marshall:1890:SAC

[34] Alfred Marshall. “Some aspects of competition.” The address of the President of Section F — Economic Science and Statistics — of the British Association, at the sixtieth meeting, held at Leeds, in September, 1890. Journal of the Royal Statistical Society, 53(4):612– 643, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979546.

Edgeworth:1890:ECCb

[35] F. Y. Edgeworth. The element of chance in competitive examinations. Journal of the Royal Statistical Society, 53(4):644– 663, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979547.

Anonymous:1890:PUS

[36] Anonymous. Population of the United States by states and territories, 1890. Journal of the Royal Statistical Society, 53(4):663–673, December 1890. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979548.

Norman:1890:SGC

[37] John Henry Norman. Silver and gold coinage of England from the conquest to the present time, &c. Journal of the Royal Statistical Society, 53(4):674– 686, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979549.

Anonymous:1890:IN

[38] Anonymous. Index numbers. Journal of the Royal Statistical Society, 53(4):686– 688, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979550.

Mouat:1890:NIC

[39] F. J. Mouat. Note on the Imperial census of India of 1891. Journal of the Royal Statistical Society, 53(4):689– 694, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979551.

Anonymous:1890:NESd

[40] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 53(4):694– 707, December 1890. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979552.

Anonymous:1890:ALd

[41] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 53(4):707–724, December 1890. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979553.

Anonymous:1891:VIa

[42] Anonymous. Volume information. Journal of the Royal Statistical Society, 54 (1):i–viii, March 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979340.

Dilke:1891:SDE

[43] Charles W. Dilke. Statistics of the defence expenditure of the Chief military and Naval powers. Journal of the Royal Statistical Society, 54(1):1–30, March 1891. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979341.

Lawson:1891:OCD

[44] Robert Lawson. The operation of the Contagious Diseases Acts among the troops in the United Kingdom, and men of the Royal Navy on the Home Station, from their introduction in 1864 to their ultimate repeal in 1884. Journal of the Royal Statistical Society, 54(1): 31–69, March 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979342.

Newsholme:1891:VSP

[45] Arthur Newsholme. The vital statistics of Peabody buildings and other artisans’ and labourers’ block dwellings. Journal of the Royal Statistical Society, 54(1): 70–111, March 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979343.

Anonymous:1891:CHR

[46] Anonymous. Commercial history and review of 1890. Journal of the Royal Statistical Society, 54(1):112–128, March 1891. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979344.

Sauerbeck:1891:PC

[47] A. Sauerbeck. Prices of commodities in 1890. Journal of the Royal Statistical Society, 54(1):128–137, March 1891. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979345.

Coste:1891:WWC

[48] Ad Coste. The wages of the working classes and the national income in France. Journal of the Royal Statistical Society, 54(1):138–153, March 1891. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979346.

Anonymous:1891:FLD

[49] Anonymous. Fires in London during the year 1890, and the Metropolitan Fire Brigade. Journal of the Royal Statistical Society, 54(1):153–158, March 1891. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979347.

Anonymous:1891:EL

[50] Anonymous. English literature in 189. Journal of the Royal Statistical Society, 54(1):158–160, March 1891. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979348.

Anonymous:1891:GL

[51] Anonymous. German literature in 1889 and 1890. Journal of the Royal Statistical Society, 54(1):160–161, March 1891. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979349.

Anonymous:1891:NESa

[52] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 54(1):161– 174, March 1891. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979350.

Anonymous:1891:ALa

[53] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 54(1):174–188, March 1891. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979351.

Anonymous:1891:PR

[54] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 54(1): 189–212, March 1891. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979352.

Mouat:1891:PEP

[55] F. J. Mouat. On prison ethics and prison labour. Journal of the Royal Statistical Society, 54(2):213–262, June 1891. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979451.

Steele:1891:CAM

[56] Dr. Steele. The charitable aspects of medical relief. Journal of the Royal Statistical Society, 54(2):263–310, June 1891. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979452.

Humphreys:1891:RRC

[57] Noel A. Humphreys. Results of the recent census and estimates of population in the largest English towns. Journal of the Royal Statistical Society, 54(2): 311–340, June 1891. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979453.

Anonymous:1891:NFP

[58] Anonymous. Note on the future provision of papers of interest and value for the consideration of the Society. Journal of the Royal Statistical Society, 54(2): 341–343, June 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979454.

Anonymous:1891:PRCa

[59] Anonymous. Preliminary return of the census in India. Journal of the Royal Statistical Society, 54(2):344–355, June 1891. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979455.

Anonymous:1891:WCP

[60] Anonymous. The white and colored population of the Southern United States, 1890. Journal of the Royal Statistical Society, 54(2):355–361, June 1891. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979456.

Anonymous:1891:EI

[61] Anonymous. Emigration and immigration in 1890. Journal of the Royal Statistical Society, 54(2):361–367, June 1891. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979457.

DeFoville:1891:ICD

[62] A. De Foville. The influence of customs duties on the price of wheat. Journal of the Royal Statistical Society, 54(2): 367–372, June 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979458.

Leroy-Beaulieu:1891:ICU

[63] P. Leroy-Beaulieu. The influence of civilisation upon the movement of the population. Journal of the Royal Statistical Society, 54(2):372–384, June 1891. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979459.

Anonymous:1891:NESb

[64] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 54(2):384–392, June 1891. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979460.

Anonymous:1891:ALb

[65] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 54(2):392–407, June 1891. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979461.

Anonymous:1891:RCF

[66] Anonymous. Report of the Council for the financial year ended 31st December, 1890, and for the sessional year ending 23rd June, 1891, presented at the fifty-seventh Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 23rd of June, 1891. Journal of the Royal Statistical Society, 54(3):409–422, September 1891. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979567.

Mouat:1891:PFS

[67] F. J. Mouat. Proceedings of the fiftyseventh Annual General Meeting. Journal of the Royal Statistical Society, 54 (3):423–428, September 1891. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979568.

Venn:1891:NUA

[68] John Venn. On the nature and uses of averages. Journal of the Royal Statistical Society, 54(3):429–456, September 1891. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979569.

Anonymous:1891:PRCb

[69] Anonymous. Preliminary report of the census of the United Kingdom. Journal of the Royal Statistical Society, 54(3): 457–491, September 1891. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979570.

Anonymous:1891:NPRa

[70] Anonymous. Notes on the preliminary returns of the censuses of some foreign countries and British colonies, 1890–91. Journal of the Royal Statistical Society, 54(3):492–502, September 1891. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979571.

Anonymous:1891:UPU

[71] Anonymous. Urban population in the United States in 1890. Journal of the Royal Statistical Society, 54(3):502– 509, September 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979572.

Nemenyi:1891:ZSH

[72] Ambroise Nemenyi. The zone system on the Hungarian railways. Journal of the Royal Statistical Society, 54(3):509– 519, September 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979573.

Morot:1891:PHF

[73] M. Ch. Morot. The progress of hippophagy in France and on the continent, as shown from statistics. Journal of the Royal Statistical Society, 54(3):519– 527, September 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979574.

Anonymous:1891:CLV

[74] Anonymous. “A century of land values:” England and Wales. Journal of the Royal Statistical Society, 54(3):528– 532, September 1891. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979575.

Anonymous:1891:PNS

[75] Anonymous. Preliminary note on the seventh International Congress of Hygiene and the sixth of Demography. Journal of the Royal Statistical Society, 54(3):532–541, September 1891. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979576.

Anonymous:1891:NESc

[76] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 54(3):542–554, September 1891. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979577.

Anonymous:1891:ALc

[77] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 54(3):555–569, September 1891. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979578.

Anonymous:1891:VIb

[78] Anonymous. Volume information. Journal of the Royal Statistical Society, 54 (4):741–754, December 1891. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979486.

Mouat:1891:OAF

[79] F. J. Mouat. The opening address of F. J. Mouat, M.D., F.R.C.S., LL.D., President of the Royal Statistical Society.

Session 1891–92. Delivered 17th November, 1891. Journal of the Royal Statistical Society, 54(4):571–596, December 1891. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979487.

Anonymous:1891:PN

[80] Anonymous. Proceedings on the 17th November, 1891. Journal of the Royal Statistical Society, 54(4):597– 599, December 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979488.

Booth:1891:ECP

[81] Charles Booth. Enumeration and classification of paupers, and state pensions for the aged. Journal of the Royal Statistical Society, 54(4):600–643, December 1891. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2979489. See

discussion [94] and correction [111].

Cunningham:1891:NCE

[82] W. Cunningham. “Nationalism and cosmopolitanism in economics.” The address of the President of the Economic Science and Statistics Section of the British Association, held at Cardiff, 1891. Journal of the Royal Statistical Society, 54(4):644–662, December 1891. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979490.

Giffen:1891:RRR

[83] R. Giffen. Report of the “Return of rates of wages in the mines and quarries in the United Kingdom”. Journal of the Royal Statistical Society, 54(4):663– 688, December 1891. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979491.

Anonymous:1891:NPRb

[84] Anonymous. Notes on the preliminary returns of the censuses of some British colonies. Journal of the Royal Statistical Society, 54(4):689–692, December 1891. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979492.

Turquan:1891:MPF

[85] V. Turquan. Movement of the population in France during the year 1890. Journal of the Royal Statistical Society, 54(4):692–696, December 1891. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979493.

Turquan:1891:CDT

[86] V. Turquan. Creation and dissolution of trades unions. Journal of the Royal Statistical Society, 54(4):697–700, December 1891. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979494.

Wagner:1891:PEP

[87] H. Wagner and A. Supan. Population of the Earth. Preface and tables. Journal of the Royal Statistical Society, 54(4):700– 710, December 1891. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979495.

Anonymous:1891:NESd

[88] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 54(4):710– 723, December 1891. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979496.

Anonymous:1891:ALd

[89] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 54(4):723–740, December 1891. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979497.

Anonymous:1892:VIa

[90] Anonymous. Volume information. Journal of the Royal Statistical Society, 55 (1):i–viii, March 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979389.

Anonymous:1892:DSJ

[91] Anonymous. Death of Sir James Caird. Journal of the Royal Statistical Society, 55(1):1, March 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979390.

Price:1892:RDA

[92] L. L. Price. The recent depression in agriculture as shown in the accounts of an Oxford College, 1876–90. Journal of the Royal Statistical Society, 55 (1):2–36, March 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979391.

Steele:1892:ADE

[93] J. C. Steele. The agricultural depression and its effects on a leading London hospital. Journal of the Royal Statistical Society, 55(1):37–55, March 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979392.

Anonymous:1892:DMB

[94] Anonymous. Discussion on Mr. Booth’s paper: Enumeration and classification of paupers and state pensions for the aged. Journal of the Royal Statistical Society, 55(1):56–79, March 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979393. See [81].

Levasseur:1892:RF

[95] E. Levasseur and R. H. H. The Russian famine. Journal of the Royal Statistical Society, 55(1):80–87, March 1892. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979394.

Baines:1892:RCI

[96] J. A. Baines. Religious census of India, 1891. Journal of the Royal Statistical Society, 55(1):88, March 1892. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979395.

Anonymous:1892:CHR

[97] Anonymous. Commercial history and review of 1891. Journal of the Royal Statistical Society, 55(1):88–109, March 1892. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979396.

Sauerbeck:1892:PC

[98] A. Sauerbeck. Prices of commodities in 1891. Journal of the Royal Statistical Society, 55(1):110–120, March 1892. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979397.

H:1892:AR

[99] R. H. H. Agricultural returns of 1891. Journal of the Royal Statistical Society, 55(1):121–126, March 1892. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979398.

H:1892:SLO

[100] R. H. H. Strikes and lock-outs in 1890. Journal of the Royal Statistical Society, 55(1):127–132, March 1892. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979399.

Anonymous:1892:P

[101] Anonymous. Pauperism in 1890–91.

Journal of the Royal Statistical Society, 55(1):132–137, March 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979400.

Anonymous:1892:FLM

[102] Anonymous. Fires in London and the Metropolitan Fire Brigade in the year 1891. Journal of the Royal Statistical Society, 55(1):137–142, March 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979401.

H:1892:FP

[103] R. H. H. Fires in Paris. Journal of the Royal Statistical Society, 55(1): 142–145, March 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979402.

H:1892:RR

[104] R. H. H. Railway returns, 1890. Journal of the Royal Statistical Society, 55(1): 145–150, March 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979403.

Anonymous:1892:EL

[105] Anonymous. English literature in 1891.

Journal of the Royal Statistical Society, 55(1):151–154, March 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979404.

Anonymous:1892:NESa

[106] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 55(1):154– 162, March 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979405.

Anonymous:1892:PR

[107] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 55(1): 163–186, March 1892. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979406.

Anonymous:1892:ALa

[108] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 55(1):187–204, March 1892. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979407.

Glover:1892:TSD

[109] John Glover. Tonnage statistics of the decade 1880–90. Journal of the Royal Statistical Society, 55(2):205–243, June 1892. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979592.

Rew:1892:ISP

[110] R. Henry Rew. An inquiry into the statistics of the production and consumption of milk and milk products in Great Britain. Journal of the Royal Statistical Society, 55(2):244–286, June 1892. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979593.

Booth:1892:EPC

[111] Charles Booth. The enumeration of paupers — a correction. Journal of the Royal Statistical Society, 55(2):287– 294, June 1892. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/

2979594. See [81].

Anonymous:1892:MRS

[112] Anonymous. The medals of the Royal Statistical Society. Journal of the Royal Statistical Society, 55(2):295–297, June 1892. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979595.

Anonymous:1892:MCK

[113] Anonymous. M. Charles Keleti. Journal of the Royal Statistical Society, 55(2): 297–298, June 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979596.

Levasseur:1892:AP

[114] Emile Levasseur. Area and population.

Journal of the Royal Statistical Society, 55(2):298–311, June 1892. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979597.

Anonymous:1892:EI

[115] Anonymous. Emigration and immigration in 1891. Journal of the Royal Statistical Society, 55(2):311–321, June 1892. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979598.

Anonymous:1892:SFU

[116] Anonymous. Sea fisheries of the United Kingdom. Journal of the Royal Statistical Society, 55(2):321–326, June 1892. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979599.

Anonymous:1892:USC

[117] Anonymous. The United States census of 1890. Journal of the Royal Statistical Society, 55(2):326–333, June 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979600.

Anonymous:1892:RCA

[118] Anonymous. Reform of the currency in Austria-Hungary. Journal of the Royal Statistical Society, 55(2):333–339, June 1892. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979601.

Anonymous:1892:NESb

[119] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 55(2):339–348, June 1892. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979602.

Anonymous:1892:ALb

[120] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 55(2):348–368, June 1892. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979603.

Anonymous:1892:RCF

[121] Anonymous. Report of the Council for the financial year ended 31st December, 1891, and for the sessional year ending 28th June, 1892, presented at the fifty-eight Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 28th of June, 1892. Journal of the Royal Statistical Society, 55(3):369–384, September 1892. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979462.

Anonymous:1892:PFE

[122] Anonymous. Proceedings of the fiftyeighth Annual General Meeting. Journal of the Royal Statistical Society, 55(3): 385–387, September 1892. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979463.

Stout:1892:SEN

[123] Robert Stout. State experiments in New Zealand. Journal of the Royal Statistical Society, 55(3):388–414, September 1892. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979464.

Fremantle:1892:APE

[124] Charles W. Fremantle. The address of the President of the Economic Science and Statistics Section of the British Association, held at Edinburgh, 1892. Journal of the Royal Statistical Society, 55(3):415–436, September 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979465.

Ogle:1892:ITO

[125] William Ogle. An inquiry into the trustworthiness of the old bills of mortality. Journal of the Royal Statistical Society, 55(3):437–460, September 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979466.

Anonymous:1892:PSF

[126] Anonymous. Proceedings of Section F of the British Association. Journal of the Royal Statistical Society, 55(3):461– 467, September 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979467.

Strachey:1892:PEL

[127] R. Strachey. On the probable effect of the limitation of the number of Ordinary Fellows elected into the Royal Society to fifteen in each year on the eventual total number of Fellows. Journal of the Royal Statistical Society, 55(3):467– 474, September 1892. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979468.

Anonymous:1892:HUS

[128] Anonymous. Homicide in the United States in 1890. Journal of the Royal Statistical Society, 55(3):474–482, September 1892. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979469.

Anonymous:1892:CIN

[129] Anonymous. A census of inland navigation in France. Journal of the Royal Statistical Society, 55(3):482–488, September 1892. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979470.

Anonymous:1892:FLN

[130] Anonymous. Female labour in New South Wales. Journal of the Royal Statistical Society, 55(3):488–492, September 1892. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979471.

Anonymous:1892:NESc

[131] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 55(3):492–506, September 1892. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979472.

Anonymous:1892:ALc

[132] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 55(3):507–520, September 1892. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979473.

Anonymous:1892:VIb

[133] Anonymous. Volume information. Journal of the Royal Statistical Society, 55 (4):709–720, December 1892. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979579.

Booth:1892:IAC

[134] Charles Booth. The inaugural address of Charles Booth, Esq., President of the Royal Statistical Society. Session 1892– 93. Delivered 15th November, 1892. Journal of the Royal Statistical Society, 55(4):521–557, December 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979580.

Anonymous:1892:PN

[135] Anonymous. Proceedings on the 15th November, 1892. Journal of the Royal Statistical Society, 55(4):558, December 1892. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979581.

Bertillion:1892:MMA

[136] Jacques Bertillion. Morbidity and mortality according to occupation. Journal of the Royal Statistical Society, 55(4): 559–600, December 1892. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979582.

Anonymous:1892:CI

[137] Anonymous. Census of Ireland, 1891.

Journal of the Royal Statistical Society, 55(4):601–619, December 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979583.

Jeans:1892:RML

[138] J. Stephen Jeans. On the recent movement of labour in different countries in reference to wages, hours of work, and efficiency. Journal of the Royal Statistical Society, 55(4):620–657, December 1892. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979584.

Mouat:1892:IPS

[139] F. J. Mouat. International prison statistics. Journal of the Royal Statistical Society, 55(4):658–668, December 1892.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979585.

Anonymous:1892:SDC

[140] Anonymous. Statistics of the damage caused by hail in Austria. Journal of the Royal Statistical Society, 55(4):668– 673, December 1892. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979586.

Anonymous:1892:PB

[141] Anonymous. Population of Bulgaria, 1888. Journal of the Royal Statistical Society, 55(4):674–675, December 1892. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979587.

Anonymous:1892:PC

[142] Anonymous. Population of Ceylon, 1891. Journal of the Royal Statistical Society, 55(4):676–677, December 1892. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979588.

Anonymous:1892:SC

[143] Anonymous. The silver census. Journal of the Royal Statistical Society, 55 (4):677–679, December 1892. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979589.

Anonymous:1892:NESd

[144] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 55(4):679– 692, December 1892. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979590.

Anonymous:1892:ALd

[145] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 55(4):693–707, December 1892. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979591.

Anonymous:1893:VIa

[146] Anonymous. Volume information. Journal of the Royal Statistical Society, 56 (1):i–viii, March 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979508.

Baines:1893:DMP

[147] J. A. Baines. Distribution and movement of the population in India. Journal of the Royal Statistical Society, 56 (1):1–43, March 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979509.

Schloss:1893:ROL

[148] David F. Schloss. The reorganisation for our labour department. Journal of the Royal Statistical Society, 56(1): 44–70, March 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979510.

Warner:1893:RIP

[149] Francis Warner. Results of an inquiry as to the physical and mental condition of fifty thousand children seen in one hundred and six schools. Journal of the Royal Statistical Society, 56(1): 71–100, March 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979511.

Anonymous:1893:CHR

[150] Anonymous. Commercial history and review of 1892. Journal of the Royal Statistical Society, 56(1):101–115, March 1893. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979512.

Anonymous:1893:AR

[151] Anonymous. Agricultural returns of 1892. Journal of the Royal Statistical Society, 56(1):116–124, March 1893. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979513.

Anonymous:1893:FLM

[152] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1892. Journal of the Royal Statistical Society, 56(1):124–129, March 1893. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979514.

Anonymous:1893:EL

[153] Anonymous. English literature in 1892. Journal of the Royal Statistical Society, 56(1):129–131, March 1893. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979515.

Anonymous:1893:POE

[154] Anonymous. Prize offered for an essay on demography. Journal of the Royal Statistical Society, 56(1):131– 132, March 1893. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979516.

Anonymous:1893:NESa

[155] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 56(1):132– 144, March 1893. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979517.

Anonymous:1893:ALa

[156] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 56(1):145–160, March 1893. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979518.

Anonymous:1893:PR

[157] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 56(1): 161–184, March 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979519.

Bourne:1893:PET

[158] Stephen Bourne. Progress of the external trade of the United Kingdom in recent years. Journal of the Royal Statistical Society, 56(2):185–214, June 1893. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979474.

Sauerbeck:1893:PCD

[159] Augustus Sauerbeck. Prices of commodities during the last seven years. Journal of the Royal Statistical Society, 56(2):215–254, June 1893. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979475.

Higgs:1893:WB

[160] Henry Higgs. Workmen’s budgets. Journal of the Royal Statistical Society, 56 (2):255–294, June 1893. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979476.

Anonymous:1893:APW

[161] Anonymous. The agricultural produce of the world. Journal of the Royal Statistical Society, 56(2):295–303, June 1893. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979477.

Anonymous:1893:LT

[162] Anonymous. Local taxation. Journal of the Royal Statistical Society, 56(2):

304–312, June 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979478.

Anonymous:1893:EI

[163] Anonymous. Emigration and immigration in 1892. Journal of the Royal Statistical Society, 56(2):312–315, June 1893. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979479.

Anonymous:1893:RGS

[164] Anonymous. Report on the Gothenburg system regulating the sale of spirituous liquor in Norway. Journal of the Royal Statistical Society, 56(2):316–327, June 1893. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979480.

Hayter:1893:DPE

[165] H. H. Hayter. Disturbance of the population estimates in Australia by defective records. Journal of the Royal Statistical Society, 56(2):327–330, June 1893. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979481.

Anonymous:1893:LDL

[166] Anonymous. The labour department and the labour gazette. Journal of the Royal Statistical Society, 56(2):330–333, June 1893. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979482.

Anonymous:1893:CPU

[167] Anonymous. Census of persons using the parks of the London County Council. Journal of the Royal Statistical Society, 56(2):333, June 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979483.

Anonymous:1893:NESb

[168] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 56(2):334–348, June 1893. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979484.

Anonymous:1893:ALb

[169] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 56(2):349–365, June 1893. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979485.

Anonymous:1893:RCF

[170] Anonymous. Report of the Council for the financial year ended 31st December, 1892, and for the sessional year ending 27th June, 1893, presented at the fifty-ninth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 27th of June, 1893. Journal of the Royal Statistical Society, 56(3):367–377, September 1893. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979634.

Anonymous:1893:PFN

[171] Anonymous. Proceedings of the fiftyninth Annual General Meeting. Journal of the Royal Statistical Society, 56(3): 378–379, September 1893. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979635.

Longstaff:1893:RD

[172] G. B. Longstaff. Rural depopulation.

Journal of the Royal Statistical Society, 56(3):380–442, September 1893. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979636.

Holmes:1893:IMF

[173] George K. Holmes. Investigations of mortgages and farm and home proprietorship in the United States. Journal of the Royal Statistical Society, 56(3):443– 481, September 1893. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979637.

Anonymous:1893:IC

[174] Anonymous. Indian currency. Journal of the Royal Statistical Society, 56(3):482– 494, September 1893. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979638.

Anonymous:1893:CSD

[175] Anonymous. Currency systems of different nations. Journal of the Royal Statistical Society, 56(3):494–505, September 1893. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979639.

Anonymous:1893:FUS

[176] Anonymous. Finances of the United States. Journal of the Royal Statistical Society, 56(3):506–518, September 1893. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979640.

Anonymous:1893:TS

[177] Anonymous. Trade of Siam. Journal of the Royal Statistical Society, 56(3):518– 522, September 1893. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979641.

Anonymous:1893:FCS

[178] Anonymous. The French “Conseil Supérieur de Statistique”. Journal of the Royal Statistical Society, 56(3):522– 524, September 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979642.

Anonymous:1893:NDF

[179] Anonymous. The national debt of France. Journal of the Royal Statistical Society, 56(3):524–525, September 1893. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979643.

Anonymous:1893:NESc

[180] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 56(3):525–541, September 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979644.

Anonymous:1893:ALc

[181] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 56(3):542–556, September 1893. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979645.

Anonymous:1893:VIb

[182] Anonymous. Volume information. Journal of the Royal Statistical Society, 56 (4):713–725, December 1893. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979430.

Booth:1893:LLP

[183] Charles Booth. Life and labour of the people in London: First results of an inquiry based on the 1891 census. Opening address of Charles Booth, Esq., President of the Royal Statistical Society. Session 1893–94. Journal of the Royal Statistical Society, 56(4):557–593, December 1893. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979431.

Anonymous:1893:PN

[184] Anonymous. Proceedings on the 21st November, 1893. Journal of the Royal Statistical Society, 56(4):594– 596, December 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979432.

deFoville:1893:WFO

[185] M. Alfred de Foville. The wealth of France and of other countries. Journal of the Royal Statistical Society, 56(4):597– 626, December 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979433.

Nicholson:1893:RFC

[186] J. S. Nicholson. The reaction in favour of the classical political economy. Address to the Economic Science and Statistical Section of the British Association, held at Nottingham, 1893. Journal of the Royal Statistical Society, 56(4):627– 639, December 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979434.

Anonymous:1893:PSF

[187] Anonymous. Proceedings of Section F of the British Association. Journal of the Royal Statistical Society, 56(4):640– 646, December 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979435.

Anonymous:1893:EFV

[188] Anonymous. Effect of the fall in value of silver on prices of commodities in China. Journal of the Royal Statistical Society, 56(4):646–669, December 1893. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979436.

Edgeworth:1893:SCB

[189] F. Y. Edgeworth. Statistical correlation between social phenomena. Journal of the Royal Statistical Society, 56(4):670– 675, December 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979437.

Pearson:1893:CMT

[190] Karl Pearson. Contributions to the mathematical theory of evolution. Journal of the Royal Statistical Society, 56 (4):675–679, December 1893. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979438.

Anonymous:1893:AI

[191] Anonymous. Alien immigration. Journal of the Royal Statistical Society, 56 (4):679–682, December 1893. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979439.

Anonymous:1893:NESd

[192] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 56(4):682– 696, December 1893. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979440.

Anonymous:1893:ALd

[193] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 56(4):696–712, December 1893. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979441.

Anonymous:1894:VIa

[194] Anonymous. Volume information. Journal of the Royal Statistical Society, 57 (1):i–vi, March 1894. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979416.

Anonymous:1894:E

[195] Anonymous. Errata. Journal of the Royal Statistical Society, 57(1):ii, March 1894. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979417.

Jones:1894:PPI

[196] Hugh R. Jones. The perils and protection of infant life. Journal of the Royal Statistical Society, 57(1):1–103, March 1894. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979418.

Hamilton:1894:OHT

[197] George Hamilton. Ocean highways: Their bearing on the food and wages of Great Britian. Journal of the Royal Statistical Society, 57(1):104–135, March 1894. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979419.

Rawson:1894:OHA

[198] Rawson W. Rawson. Ocean highways: Approaches to the United Kingdom. Journal of the Royal Statistical Society, 57(1):136–148, March 1894. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979420.

Anonymous:1894:CHR

[199] Anonymous. Commercial history and review of 1893. Journal of the Royal Statistical Society, 57(1):149–165, March 1894. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979421.

Anonymous:1894:ARa

[200] Anonymous. Agricultural returns of 1893. Journal of the Royal Statistical Society, 57(1):165–168, March 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979422.

Anonymous:1894:ISI

[201] Anonymous. The International Statistical Institute at Chicago. Journal of the Royal Statistical Society, 57(1): 168–171, March 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979423.

Sauerbeck:1894:PC

[202] A. Sauerbeck. Prices of commodities in 1893. Journal of the Royal Statistical Society, 57(1):172–183, March 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979424.

Anonymous:1894:FLM

[203] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1893. Journal of the Royal Statistical Society, 57(1):184–189, March 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979425.

Anonymous:1894:EL

[204] Anonymous. English literature in 1893.

Journal of the Royal Statistical Society, 57(1):189–192, March 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979426.

Anonymous:1894:SEA

[205] Anonymous. Statistical and economical article in recent periodicals. Journal of the Royal Statistical Society, 57(1): 192–196, March 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979427.

Anonymous:1894:ALa

[206] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 57(1):196–211, March 1894. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979428.

Anonymous:1894:PR

[207] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 57(1): 212–234, March 1894. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979429.

Booth:1894:SPO

[208] Charles Booth. Statistics of pauperism in old age. Journal of the Royal Statistical Society, 57(2):235–253, June 1894. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979498.

Baines:1894:CPP

[209] J. A. Baines. Conditions and prospects of popular education in India. Journal of the Royal Statistical Society, 57(2): 254–297, June 1894. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979499.

Hooker:1894:MCT

[210] Reginald H. Hooker. Modes of censustaking in the British Dominions. Journal of the Royal Statistical Society, 57(2): 298–368, June 1894. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979500.

Anonymous:1894:TPG

[211] Anonymous. Tables of the production of gold and silver in the world since the discovery of America. Journal of the Royal Statistical Society, 57(2):369–378, June 1894. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979501.

Anonymous:1894:CEW

[212] Anonymous. Census of England and Wales: Deputation to the President of the local government board. Journal of the Royal Statistical Society, 57(2): 379–383, June 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979502.

Anonymous:1894:ARb

[213] Anonymous. Agricultural returns of 1893. Journal of the Royal Statistical Society, 57(2):383–393, June 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979503.

Bateman:1894:CTS

[214] A. E. Bateman. Comparability of trade statistics of various countries. Journal of the Royal Statistical Society, 57(2): 394–399, June 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979504.

Anonymous:1894:SCT

[215] Anonymous. Statistics of the consumption of tea. Journal of the Royal Statistical Society, 57(2):399–400, June 1894. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979505.

Anonymous:1894:NESa

[216] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 57(2):401–417, June 1894. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979506.

Anonymous:1894:ALb

[217] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 57(2):418–436, June 1894. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979507.

Anonymous:1894:RCF

[218] Anonymous. Report of the Council for the financial year ended 31st December, 1893, and for the sessional year ending 26th June, 1894, presented at the sixtieth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 26th of June, 1894. Journal of the Royal Statistical Society, 57(3):437–449, September 1894. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979324.

Anonymous:1894:PSA

[219] Anonymous. Proceedings of the sixtieth Annual General Meeting. Journal of the Royal Statistical Society, 57(3):450– 451, September 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979325.

Macdonell:1894:SLE

[220] John Macdonell. Statistics of litigation in England and Wales since 1859. Journal of the Royal Statistical Society, 57(3):452–523, September 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979326.

Harris:1894:CGW

[221] William J. Harris. A comparison of the growth of wealth in France and England, also of their economic conditions, specially with reference to their agricultural systems, and their position in case of war. Journal of the Royal Statistical Society, 57(3):524–562, September 1894. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979327.

Edgeworth:1894:ACB

[222] F. Y. Edgeworth. Asymmetrical correlation between social phenomena. Journal of the Royal Statistical Society, 57(3): 563–568, September 1894. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979328.

Anonymous:1894:ARG

[223] Anonymous. Agricultural returns of Great Britain, 1894. Journal of the Royal Statistical Society, 57(3):568–569, September 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979329.

Anonymous:1894:NESb

[224] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 57(3):570–579, September 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979330.

Anonymous:1894:ALc

[225] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 57(3):579–594, September 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979331.

Anonymous:1894:VIb

[226] Anonymous. Volume information. Journal of the Royal Statistical Society, 57 (4):749–760, December 1894. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979604.

Farrer:1894:RBM

[227] Lord Farrer. The relations between morals, economics, and statistics. inaugural address of the Right Hon. Lord Farrer, President of the Royal Statistical Society, session 1894–95, delivered 20th November, 1894. Journal of the Royal Statistical Society, 57(4):595– 608, December 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979605.

Anonymous:1894:PN

[228] Anonymous. Proceedings on the 20th November, 1894. Journal of the Royal Statistical Society, 57(4):609– 610, December 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979606.

Bastable:1894:AES

[229] C. F. Bastable. Address to the economic science and statistics section of the British Association, held at Oxford, 1894. Journal of the Royal Statistical Society, 57(4):611–626, December 1894. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979607.

Hooker:1894:RBW

[230] R. H. Hooker. On the relation between wages and the numbers employed in the coal mining industry. Journal of the Royal Statistical Society, 57(4):627– 642, December 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979608.

Porter:1894:EUS

[231] Robert P. Porter. The eleventh United States census. Journal of the Royal Statistical Society, 57(4):643–677, December 1894. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979609.

Hollerith:1894:ETM

[232] Herman Hollerith. The electrical tabulating machine. Journal of the Royal Statistical Society, 57(4):678– 689, December 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979610.

Korosi:1894:EDL

[233] Joseph Korosi. An estimate of the degree of legitimate natality drawn from observations made at Budapest. Journal of the Royal Statistical Society, 57(4):690– 702, December 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979611.

Galton:1894:RDN

[234] Francis Galton. “Results derived from the natality table of Korosi by employing the method of contours or isogens”. Journal of the Royal Statistical Society, 57(4):702–708, December 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979612.

Anonymous:1894:PSF

[235] Anonymous. Proceedings of Section F of the British Association. Journal of the Royal Statistical Society, 57(4):708– 714, December 1894. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979613.

Fox:1894:STC

[236] C. Fox. Statistics of tea consumption.

Journal of the Royal Statistical Society, 57(4):714–716, December 1894. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979614.

Anonymous:1894:CP

[237] Anonymous. Census of Portugal, 1890.

Journal of the Royal Statistical Society, 57(4):717, December 1894. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979615.

Anonymous:1894:NESc

[238] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 57(4):717– 732, December 1894. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979616.

Anonymous:1894:ALd

[239] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 57(4):732–748, December 1894. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979617.

Anonymous:1895:VIa

[240] Anonymous. Volume information. Journal of the Royal Statistical Society, 58 (1):i–vi, March 1895. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979554.

Drage:1895:AI

[241] Geoffrey Drage. Alien immigration. Journal of the Royal Statistical Society, 58(1):1–35, March 1895. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979555.

Price:1895:COA

[242] L. L. Price. The colleges of Oxford and agricultural depression. Journal of the Royal Statistical Society, 58(1): 36–74, March 1895. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979556.

Crawford:1895:IWP

[243] R. F. Crawford. An inquiry into wheat prices and wheat supply. Journal of the Royal Statistical Society, 58(1):75– 120, March 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979557.

Anonymous:1895:CHR

[244] Anonymous. Commercial history and review of 1894. Journal of the Royal Statistical Society, 58(1):121–139, March 1895. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979558.

Anonymous:1895:PC

[245] Anonymous. Prices of commodities in 1894. Journal of the Royal Statistical Society, 58(1):140–154, March 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979559.

Anonymous:1895:FRB

[246] Anonymous. Financial relations between Great Britain and Ireland. Journal of the Royal Statistical Society, 58(1): 155–163, March 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979560.

Anonymous:1895:BH

[247] Anonymous. The British harvest of 1894. Journal of the Royal Statistical Society, 58(1):163–164, March 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979561.

Anonymous:1895:FLM

[248] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1894. Journal of the Royal Statistical Society, 58(1):164–170, March 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979562.

Anonymous:1895:EL

[249] Anonymous. English literature in 1894.

Journal of the Royal Statistical Society, 58(1):171–172, March 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979563.

Anonymous:1895:NESa

[250] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 58(1):173– 184, March 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979564.

Anonymous:1895:QLAa

[251] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 58(1):184–199, March 1895. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979565.

Anonymous:1895:PR

[252] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 58(1): 200–222, March 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979566.

Bowley:1895:CAW

[253] A. L. Bowley. Changes in average wages (nominal and real) in the United Kingdom between 1860 and 1891. Journal of the Royal Statistical Society, 58(2): 223–285, June 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979353.

Brabrook:1895:PFS

[254] E. W. Brabrook. The progress of friendly societies and other institutions connected with the friendly societies registry Office during the ten years 1884– 94. Journal of the Royal Statistical Society, 58(2):286–302, June 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979354.

Wilkinson:1895:SIF

[255] J. Frome Wilkinson. Some illustrations of Friendly Society finance. Journal of the Royal Statistical Society, 58(2): 303–326, June 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979355.

Smith:1895:MFL

[256] Edward Orford Smith. Municipal finance, or local taxation and local expenditure, as illustrated by the case of the City of Birmingham. Journal of the Royal Statistical Society, 58(2):327–364, June 1895. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979356.

Anonymous:1895:CP

[257] Anonymous. Corn prices, 1806–61.

Journal of the Royal Statistical Society, 58(2):365–367, June 1895. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979357.

Anonymous:1895:RSD

[258] Anonymous. Remarks on and suggestions for a decimal currency. Journal of the Royal Statistical Society, 58(2): 367–369, June 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979358.

Anonymous:1895:NESb

[259] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 58(2):370–383, June 1895. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979359.

Anonymous:1895:QLAb

[260] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 58(2):384–400, June 1895. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979360.

Anonymous:1895:RCF

[261] Anonymous. Report of the Council for the financial year ended 31st December, 1894, and for the sessional year ending 25th June, 1895, presented at the sixty-first Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 25th of June, 1895. Journal of the Royal Statistical Society, 58(3):401–412, September 1895. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979532.

Farrer:1895:PSF

[262] Lord Farrer. Proceedings of the sixtyfirst Annual General Meeting. Journal of the Royal Statistical Society, 58(3):413– 416, September 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979533.

Robertson:1895:SSB

[263] J. Barr Robertson. Some statistics bearing upon bimetallism. Journal of the Royal Statistical Society, 58(3):417– 479, September 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979534.

Nowell:1895:CSP

[264] Edwin C. Nowell. Comparative statistics of the principal British possessions and foreign countries. Journal of the Royal Statistical Society, 58(3):480– 505, September 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979535.

Edgeworth:1895:SRC

[265] F. Y. Edgeworth. On some recent contributions to the theory of statistics. Journal of the Royal Statistical Society, 58(3):506–515, September 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979536.

Anonymous:1895:DPI

[266] Anonymous. Distribution of property and incomes in Norway. Journal of the Royal Statistical Society, 58(3):515– 518, September 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979537.

Anonymous:1895:SEW

[267] Anonymous. Statistics of employment of women and girls. Journal of the Royal Statistical Society, 58(3):518–528, September 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979538.

Anonymous:1895:SBE

[268] Anonymous. Statistics of bankruptcies in England and Wales, 1885–94. Journal of the Royal Statistical Society, 58(3): 528–530, September 1895. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979539.

Anonymous:1895:NESc

[269] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 58(3):531–542, September 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979540.

Anonymous:1895:QLAc

[270] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 58(3):542–558, September 1895. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979541.

Anonymous:1895:VIb

[271] Anonymous. Volume information. Journal of the Royal Statistical Society, 58 (4):675–686, December 1895. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979674.

Probyn:1895:GSM

[272] Lesley C. Probyn. Gold and silver and the money of the world. Journal of the Royal Statistical Society, 58(4):559– 590, December 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979675.

Price:1895:RES

[273] L. L. Price. On the relations of economic science to practical affairs. Address to the economic science and statistics section of the British Association, held at Ipswich, 1895. Journal of the Royal Statistical Society, 58(4):591–609, December 1895. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979676.

Anonymous:1895:ITI

[274] Anonymous. Incidence of taxation on the income of an individual in France. Journal of the Royal Statistical Society, 58(4):610–625, December 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979677.

Anonymous:1895:RAB

[275] Anonymous. Raiffeisen (agricultural) banks in Germany, Italy, and Belgium. Journal of the Royal Statistical Society, 58(4):626–631, December 1895. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979678.

Anonymous:1895:ISI

[276] Anonymous. The International Statistical Institute. Journal of the Royal Statistical Society, 58(4):631–635, December 1895. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979679.

Anonymous:1895:GCI

[277] Anonymous. A general census in 1900.

inquiry as to the possibility of making it synchronous, and as to the extent to which it can be made uniform. Journal of the Royal Statistical Society, 58 (4):635–638, December 1895. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979680.

Anonymous:1895:ARG

[278] Anonymous. Agricultural returns of Great Britain, 1895. Journal of the Royal Statistical Society, 58(4):638– 641, December 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979681.

Anonymous:1895:NESd

[279] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 58(4):642– 658, December 1895. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979682.

Anonymous:1895:QLAd

[280] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 58(4):659–674, December 1895. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979683.

Anonymous:1896:VIa

[281] Anonymous. Volume information. Journal of the Royal Statistical Society, 59 (1):i–vi, March 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979752.

Newsholme:1896:NSN

[282] Arthur Newsholme. A national system of notification and registration of sickness. Journal of the Royal Statistical Society, 59(1):1–37, March 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979753.

Baines:1896:PRE

[283] J. A. Baines. Parliamentary representation in England illustrated by the elections of 1892 and 1895. Journal of the Royal Statistical Society, 59(1): 38–124, March 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979754.

Warner:1896:MPC

[284] Francis Warner. Mental and physical conditions among fifty thousand children seen 1892–94 seen 1892–94, and the methods of studing recorded observations, with special reference to the determination of the causes of mental dulness and other defects. Journal of the Royal Statistical Society, 59(1): 125–168, March 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979755.

Anonymous:1896:CHR

[285] Anonymous. Commercial history and review of 1895. Journal of the Royal Statistical Society, 59(1):169–186, March 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979756.

Sauerbeck:1896:PC

[286] A. Sauerbeck. Prices of commodities in 1895. Journal of the Royal Statistical Society, 59(1):186–201, March 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979757.

Anonymous:1896:FLM

[287] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1895. Journal of the Royal Statistical Society, 59(1):202–207, March 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979758.

Anonymous:1896:EL

[288] Anonymous. English literature in 1895.

Journal of the Royal Statistical Society, 59(1):207–209, March 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979759.

Anonymous:1896:SEA

[289] Anonymous. Statistical and economical articles in recent periodicals. Journal of the Royal Statistical Society, 59(1): 209–214, March 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979760.

Anonymous:1896:QLAa

[290] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 59(1):214–231, March 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979761.

Anonymous:1896:PR

[291] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 59(1): 232–254, March 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979762.

Watson:1896:RIS

[292] John Watson. Reformatory and industrial schools. Journal of the Royal Statistical Society, 59(2):255–317, June 1896. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979776.

Yule:1896:NHP

[293] G. Udny Yule. Notes on the history of pauperism in England and Wales from 1850, treated by the method of frequency-curves; with an introduction on the method. Journal of the Royal Statistical Society, 59(2):318–357, June 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979777.

Edgeworth:1896:SUC

[294] F. Y. Edgeworth. Statistics of unprogressive communities. Journal of the Royal Statistical Society, 59(2):358–386, June 1896. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979778.

Longstaff:1896:PPU

[295] G. B. Longstaff. The present population of the United Kingdom, 5th April, 1896. Journal of the Royal Statistical Society, 59(2):387–392, June 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979779.

Galton:1896:AMP

[296] Francis Galton. Application of the method of percentiles to Mr. Yule’s data on the distribution of pauperism. Journal of the Royal Statistical Society, 59 (2):392–396, June 1896. ISSN 0952-

8385. URL http://www.jstor.org/

stable/2979780. See remarks [297].

Yule:1896:RMG

[297] G. U. Yule. Remarks on Mr. Galton’s note. Journal of the Royal Statistical Society, 59(2):396–398, June 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979781. See [296].

Pearson:1896:CMT

[298] Karl Pearson and F. Galton. “Contributions to the mathematical theory of evolution. Note on reproductive selection.”. Journal of the Royal Statistical Society, 59(2):398–402, June 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979782.

Anonymous:1896:BA

[299] Anonymous. British agriculture in 1895.

Journal of the Royal Statistical Society, 59(2):402–406, June 1896. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979783.

Anonymous:1896:NESa

[300] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 59(2):407–426, June 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979784.

Anonymous:1896:QLAb

[301] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 59(2):426–442, June 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979785.

Anonymous:1896:RCF

[302] Anonymous. Report of the Council for the financial year ending 31st December, 1895, and for the sessional year during ending 26th June, 1896, presented at the sixty-second Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 26 June, 1896. Journal of the Royal Statistical Society, 59(3):443–454, September 1896. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979826.

Rawson:1896:PSS

[303] Rawson W. Rawson. Proceedings of the sixty-second Annual General Meeting. Journal of the Royal Statistical Society, 59(3):455–458, September 1896. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979827.

Yerburgh:1896:ACB

[304] R. A. Yerburgh. Agricultural credit banks. Journal of the Royal Statistical Society, 59(3):459–484, September 1896. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979828.

Price-Williams:1896:RRT

[305] R. Price-Williams. Railway rates and terminal charges. Journal of the Royal Statistical Society, 59(3):485–528, September 1896. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979829.

Edgeworth:1896:SNS

[306] F. Y. Edgeworth. Supplementary notes on statistics. Journal of the Royal Statistical Society, 59(3):529–539, September 1896. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979830.

Neymarck:1896:DPP

[307] A. Neymarck. The distribution of personal property in France. Journal of the Royal Statistical Society, 59(3):540– 543, September 1896. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979831.

Humphreys:1896:RGD

[308] Noel A. Humphreys. The Registrar-General’s decennial supplement, 1881– 90. Journal of the Royal Statistical Society, 59(3):543–546, September 1896. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979832.

Anonymous:1896:LC

[309] Anonymous. The London census, 1896.

Journal of the Royal Statistical Society, 59(3):546–548, September 1896. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979833.

Anonymous:1896:ARGa

[310] Anonymous. Agricultural returns of Great Britain, 1896. Journal of the Royal Statistical Society, 59(3):548–549, September 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979834.

Anonymous:1896:NESb

[311] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 59(3):550–562, September 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979835.

Anonymous:1896:QLAc

[312] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 59(3):563–578, September 1896. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979836.

Anonymous:1896:VIb

[313] Anonymous. Volume information. Journal of the Royal Statistical Society, 59 (4):709–720, December 1896. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979898.

Martin:1896:SDS

[314] John Biddulph Martin. On some developments of statistical research and methods during recent years. The inaugural address of John Biddulph Martin, Esq., M.A., President of the Royal Statistical Society. Session 1896–97. Delivered 17th November, 1896. Journal of the Royal Statistical Society, 59(4):579– 628, December 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979899.

Anonymous:1896:PN

[315] Anonymous. Proceedings on the 17th November, 1896. Journal of the Royal Statistical Society, 59(4):629– 630, December 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979900.

Courtney:1896:AES

[316] Leonard Courtney. Address to the economic science and statistics section of the British Association, held at Liverpool, 1896. Journal of the Royal Statistical Society, 59(4):631–643, December 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979901.

Blunden:1896:SOD

[317] G. H. Blunden. Some observations on the distribution and incidence of rates and taxes; with special reference to the transfer of charges from the former to the latter. Journal of the Royal Statistical Society, 59(4):644–656, December 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979902.

Baines:1896:TI

[318] J. A. Baines. The trade of India in 1895–96. Journal of the Royal Statistical Society, 59(4):657–664, December 1896. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979903.

Anonymous:1896:COG

[319] Anonymous. Census of occupations in Germany, 1895. Journal of the Royal Statistical Society, 59(4):664– 669, December 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979904.

Anonymous:1896:DCT

[320] Anonymous. The depression in the coal trade. Journal of the Royal Statistical Society, 59(4):669–673, December 1896. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979905.

Anonymous:1896:ARGb

[321] Anonymous. Agricultural returns of Great Britain for 1896. Journal of the Royal Statistical Society, 59(4):673– 675, December 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979906.

Anonymous:1896:NESc

[322] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 59(4):676– 692, December 1896. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979907.

Anonymous:1896:QLAd

[323] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 59(4):692–708, December 1896. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979908.

Anonymous:1897:VIa

[324] Anonymous. Volume information. Journal of the Royal Statistical Society, 60 (1):i–vi, March 1897. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979684.

Morrison:1897:ICS

[325] William Douglas Morrison. The interpretation of criminal statistics. Journal of the Royal Statistical Society, 60 (1):1–32, March 1897. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979685.

Welton:1897:LDR

[326] Thomas A. Welton. Local death-rates in England and Wales in the ten years 1881–90. Journal of the Royal Statistical Society, 60(1):33–83, March 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979686.

Atkinson:1897:SPI

[327] Fred J. Atkinson. Silver prices in India. Journal of the Royal Statistical Society, 60(1):84–147, March 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979687.

Rubin:1897:MC

[328] Marcus Rubin and A. W. Flux. A measure of civilisation. Journal of the Royal Statistical Society, 60(1):148– 161, March 1897. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979688.

Anonymous:1897:RCR

[329] Anonymous. Report of a committee with reference to the proposed universal synchronous census in 1900. Journal of the Royal Statistical Society, 60(1): 162–163, March 1897. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979689.

Anonymous:1897:CHR

[330] Anonymous. Commercial history and review of 1896. Journal of the Royal Statistical Society, 60(1):164–179, March 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979690.

Sauerbeck:1897:PC

[331] A. Sauerbeck. Prices of commodities in 1896. Journal of the Royal Statistical Society, 60(1):180–194, March 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979691.

Anonymous:1897:FLM

[332] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1896. Journal of the Royal Statistical Society, 60(1):195–198, March 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979692.

Anonymous:1897:EL

[333] Anonymous. English literature in 1896.

Journal of the Royal Statistical Society, 60(1):198–200, March 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979693.

Anonymous:1897:NESa

[334] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 60(1):200– 211, March 1897. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979694.

Anonymous:1897:QLAa

[335] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 60(1):212–227, March 1897. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979695.

Anonymous:1897:PR

[336] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 60(1): 228–250, March 1897. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979696.

Danvers:1897:AED

[337] Frederick Charles Danvers. Agriculture in Essex during the past fifty years, as exemplified by the records of one farm, with special reference to the prices of corn and the conditions of labour. Journal of the Royal Statistical Society, 60 (2):251–277, June 1897. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979799.

Wolff:1897:SBH

[338] Henry W. Wolff. Savings banks at home and abroad. Journal of the Royal Statistical Society, 60(2):278–359, June 1897. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979800.

Hunter:1897:MIP

[339] Robert Hunter. The movements for the inclosure and preservation of open lands. Journal of the Royal Statistical Society, 60(2):360–431, June 1897. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979801.

Anonymous:1897:OMJ

[340] Anonymous. [obituary]: Mr. John Biddulph Martin, M.A. Journal of the Royal Statistical Society, 60(2):432–434, June 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979802.

Anonymous:1897:ODF

[341] Anonymous. [obituary]: Dr. Frederic John Mouat, M.D., LL.D. Journal of the Royal Statistical Society, 60(2):434–437, June 1897. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979803.

Bowley:1897:CPI

[342] A. L. Bowley. Changes of prices of imports and exports since 1881. Journal of the Royal Statistical Society, 60(2): 437–439, June 1897. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979804.

Pearson:1897:MCT

[343] Karl Pearson and Alice Lee. Mathematical contributions to the theory of evolution. On telegony in man, &c. “. Journal of the Royal Statistical Society, 60(2): 440–449, June 1897. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979805.

Anonymous:1897:AR

[344] Anonymous. The agricultural returns of 1896. Journal of the Royal Statistical Society, 60(2):449–453, June 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979806.

Anonymous:1897:NSA

[345] Anonymous. Number and size of agricultural holdings. Journal of the Royal Statistical Society, 60(2):453–455, June 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979807.

Anonymous:1897:VRF

[346] Anonymous. The Victoria research fund. Journal of the Royal Statistical Society, 60(2):455–456, June 1897. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979808.

Yule:1897:NTT

[347] G. Udny Yule. Note on the teaching of the theory of statistics at University College. Journal of the Royal Statistical Society, 60(2):456–458, June 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979809.

Anonymous:1897:NESb

[348] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 60(2):458–472, June 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979810.

Anonymous:1897:QLAb

[349] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 60(2):472–486, June 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979811.

Anonymous:1897:RCF

[350] Anonymous. Report of the Council for the financial year ended 31st December, 1896, and for the sessional year ending 29th June, 1897, presented at the sixty-third Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 29th of June, 1897. Journal of the Royal Statistical Society, 60(3):487–499, September 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979709.

Bateman:1897:PST

[351] A. E. Bateman. Proceedings of the sixtythird Annual General Meeting. Journal of the Royal Statistical Society, 60(3): 500–502, September 1897. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979710.

Humphreys:1897:EVS

[352] Noel A. Humphreys. English vaccination and small-pox statistics; with special reference to the report of the Royal Commission, and to recent small-pox epidemics. Journal of the Royal Statistical Society, 60(3):503–551, September 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979711.

Milnes:1897:SSP

[353] Alfred Milnes. Statistics of small-pox and vaccination, with special reference to age-incidence, sex-incidence, and sanitation. Journal of the Royal Statistical Society, 60(3):552–612, September 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979712.

Kerr:1897:SHM

[354] James Kerr. School hygiene, in its mental, moral, and physical aspects. Journal of the Royal Statistical Society, 60(3): 613–680, September 1897. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979713.

Edgeworth:1897:MAC

[355] F. Y. Edgeworth. Miscellaneous applications of the calculus of probabilities. Journal of the Royal Statistical Society, 60(3):681–698, September 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979714.

Sheppard:1897:CAS

[356] W. F. Sheppard. On the calculation of the average square, cube, of a large number of magnitudes. Journal of the Royal Statistical Society, 60(3):698– 703, September 1897. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979715.

Anonymous:1897:RRC

[357] Anonymous. The report of the Royal Commission on Agriculture. Journal of the Royal Statistical Society, 60(3):704– 707, September 1897. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979716.

Anonymous:1897:ARGa

[358] Anonymous. Agricultural returns of Great Britain, 1897. Journal of the Royal Statistical Society, 60(3):707–708, September 1897. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979717.

Anonymous:1897:NESc

[359] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 60(3):709–721, September 1897. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979718.

Anonymous:1897:QLAc

[360] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 60(3):722–734, September 1897. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979719.

Anonymous:1897:VIb

[361] Anonymous. Volume information. Journal of the Royal Statistical Society, 60 (4):899–912, December 1897. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979742.

Craigie:1897:NSD

[362] P. G. Craigie. Notes on the subjects discussed at the St. Petersburg meeting of the International Statistical Institute. Journal of the Royal Statistical Society, 60(4):735–788, December 1897. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979743.

Courtney:1897:JCQ

[363] Leonard H. Courtney. “Jevons’s coal question: Thirty years after.” The inaugural address of the Right Hon. Leonard H. Courtney, M.A., M.P., President of the Royal Statistical Society. Session 1897–98. Delivered 14th December, 1897. Journal of the Royal Statistical Society, 60(4):789–810, December 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979744.

Anonymous:1897:PD

[364] Anonymous. Proceedings on the 14th December, 1897. Journal of the Royal Statistical Society, 60(4):811, December 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979745.

Yule:1897:TC

[365] G. Udny Yule. On the theory of correlation. Journal of the Royal Statistical Society, 60(4):812–854, December 1897. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979746.

Bowley:1897:RBA

[366] A. L. Bowley. Relations between the accuracy of an average and that of its constituent parts. Journal of the Royal Statistical Society, 60(4):855–866, December 1897. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979747.

Flux:1897:CCM

[367] A. W. Flux. The cost of compensation for mining accidents in Germany. Journal of the Royal Statistical Society, 60(4):866–871, December 1897. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979748.

Anonymous:1897:ARGb

[368] Anonymous. Agricultural returns of Great Britain for 1897. Journal of the Royal Statistical Society, 60(4):872– 874, December 1897. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979749.

Anonymous:1897:NESd

[369] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 60(4):874– 883, December 1897. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979750.

Anonymous:1897:QLAd

[370] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 60(4):883–898, December 1897. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979751.

Anonymous:1898:VIa

[371] Anonymous. Volume information. Journal of the Royal Statistical Society, 61 (1):i–vi, March 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979786.

Baines:1898:RCT

[372] J. A. Baines. The recent course of trade within the British empire. Journal of the Royal Statistical Society, 61 (1):1–48, March 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979787.

Cannan:1898:DSU

[373] Edwin Cannan. Demographic statistics of the United Kingdom: Their want of correlation and other defects. Journal of the Royal Statistical Society, 61(1): 49–82, March 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979788.

Atkinson:1898:SAR

[374] Fred J. Atkinson. The statistical aspects of the recent bimetallic proposals. Journal of the Royal Statistical Society, 61(1):83–117, March 1898. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979789.

Heywood:1898:JHM

[375] James Heywood. James Heywood, M.A., F.R.S., F.G.S. Journal of the Royal Statistical Society, 61(1):118–119, March 1898. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979790.

Edgeworth:1898:MACa

[376] F. Y. Edgeworth. Miscellaneous applications of the calculus of probabilities. Journal of the Royal Statistical Society, 61(1):119–131, March 1898. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979791.

Anonymous:1898:CHR

[377] Anonymous. Commercial history and review of 1897. Journal of the Royal Statistical Society, 61(1):131–148, March 1898. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979792.

Sauerbeck:1898:PC

[378] A. Sauerbeck. Prices of commodities in 1897. Journal of the Royal Statistical Society, 61(1):149–162, March 1898. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979793.

Anonymous:1898:FLM

[379] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1897. Journal of the Royal Statistical Society, 61(1):163–166, March 1898. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979794.

Anonymous:1898:EL

[380] Anonymous. English literature in 1897. Journal of the Royal Statistical Society, 61(1):167–168, March 1898. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979795.

Anonymous:1898:NESa

[381] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 61(1):169– 180, March 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979796.

Anonymous:1898:QLAa

[382] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 61(1):180–195, March 1898. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979797.

Anonymous:1898:PR

[383] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 61(1): 196–218, March 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979798.

Collet:1898:CUO

[384] Clara E. Collet. The collection and utilisation of official statistics bearing on the extent and effects of the industrial employment of women. Journal of the Royal Statistical Society, 61(2):219–270, June 1898. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979733.

Loch:1898:PRS

[385] C. S. Loch. Poor relief in Scotland:

Its statistics and development, 1791 to 1891. Journal of the Royal Statistical Society, 61(2):271–370, June 1898. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979734.

Anonymous:1898:RHW

[386] Anonymous. The Right Hon. W. E.

Gladstone. Journal of the Royal Statistical Society, 61(2):371–372, June 1898. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979735.

Anonymous:1898:JTD

[387] Anonymous. John Towne Danson. Journal of the Royal Statistical Society, 61 (2):372–374, June 1898. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979736.

Baines:1898:CMD

[388] J. A. Baines. On census matters discussed at the St. Petersburg meeting of the International Statistical Institute, 1897. Journal of the Royal Statistical Society, 61(2):374–382, June 1898. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979737.

Anonymous:1898:AR

[389] Anonymous. The agricultural returns of 1897. Journal of the Royal Statistical Society, 61(2):382–387, June 1898. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979738.

Atkinson:1898:SPI

[390] Fred J. Atkinson. Silver prices in India for 1897. Journal of the Royal Statistical Society, 61(2):387–403, June 1898. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979739.

Anonymous:1898:NESb

[391] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 61(2):403–412, June 1898. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979740.

Anonymous:1898:QLAb

[392] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 61(2):412–428, June 1898. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979741.

Anonymous:1898:RCF

[393] Anonymous. Report of the Council for the financial year ended 31st December, 1897, and for the sessional year ending 28th June, 1898, presented at the sixty-fourth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 28th of June, 1898. Journal of the Royal Statistical Society, 61(3):429–440, September 1898. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979646.

Anonymous:1898:PSF

[394] Anonymous. Proceedings of the sixtyfourth Annual General Meeting. Journal of the Royal Statistical Society, 61(3): 441, September 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979647.

Gomme:1898:LTL

[395] Laurence Gomme. Local taxation in London. Journal of the Royal Statistical Society, 61(3):442–517, September 1898. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979648.

Coghlan:1898:DCB

[396] T. A. Coghlan. Deaths in child-birth in New South Wales. Journal of the Royal Statistical Society, 61(3):518– 533, September 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2979649. See reply [420].

Edgeworth:1898:MACb

[397] F. Y. Edgeworth. Miscellaneous applications of the calculus of probabilities. Journal of the Royal Statistical Society, 61(3):534–544, September 1898. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979650.

Baines:1898:LSU

[398] J. A. Baines. London’s “storied urn,”

the statistical aspect of recent County Council elections. Journal of the Royal Statistical Society, 61(3):544–560, September 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979651.

Anonymous:1898:IMC

[399] Anonymous. Increased mortality from cancer. Journal of the Royal Statistical Society, 61(3):560–564, September 1898. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979652.

Anonymous:1898:ARGa

[400] Anonymous. Agricultural returns of Great Britain, 1898. Journal of the Royal Statistical Society, 61(3):564–565, September 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979653.

Anonymous:1898:NESc

[401] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 61(3):565–580, September 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979654.

Anonymous:1898:QLAc

[402] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 61(3):581–596, September 1898. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979655.

Anonymous:1898:VIb

[403] Anonymous. Volume information. Journal of the Royal Statistical Society, 61 (4):757–771, December 1898. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979850.

Burdett:1898:OAP

[404] Henry Burdett. Old age pensions. Journal of the Royal Statistical Society, 61 (4):597–639, December 1898. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979851.

Anonymous:1898:ECE

[405] Anonymous. “An experiment in commercial expansion.” The annual address of the Right Hon. Leonard H. Courtney, M.A., M.P., President of the Royal Statistical Society. Session 1898–99. Delivered 13th December, 1898. Journal of the Royal Statistical Society, 61(4):640– 667, December 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979852.

Anonymous:1898:PD

[406] Anonymous. Proceedings on the 13th December, 1898. Journal of the Royal Statistical Society, 61(4):668– 669, December 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979853.

Edgeworth:1898:RSM

[407] F. Y. Edgeworth. On the representation of statistics by mathematical formulæ (part i.). Journal of the Royal Statistical Society, 61(4):670–700, December 1898. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979854.

Salmon:1898:DCB

[408] R. G. Salmon. Deaths in child-birth; with a table showing the probable number of children from marriage between couples at certain ages. Journal of the Royal Statistical Society, 61(4):701– 702, December 1898. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979855.

Bowley:1898:SWU

[409] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (part i.) agricultural wages. Journal of the Royal Statistical Society, 61(4):702–722, December 1898. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979856.

Tupp:1898:LIC

[410] A. Cotterell Tupp. Longevity in the Indian civil service. Journal of the Royal Statistical Society, 61(4):722– 725, December 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979857.

Anonymous:1898:ARGb

[411] Anonymous. Agricultural returns of Great Britain for 1898. Journal of the Royal Statistical Society, 61(4):725– 727, December 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979858.

Anonymous:1898:NESd

[412] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 61(4):728– 740, December 1898. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979859.

Anonymous:1898:QLAd

[413] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 61(4):740–756, December 1898. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979860.

Anonymous:1899:VIa

[414] Anonymous. Volume information. Journal of the Royal Statistical Society, 62 (1):i–viii, March 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979875.

Giffen:1899:EI

[415] Robert Giffen. The excess of imports.

Journal of the Royal Statistical Society, 62(1):1–82, March 1899. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979876.

Howell:1899:CSA

[416] Price Howell. Comparative statistics of Australasian railways. Journal of the Royal Statistical Society, 62(1):83– 124, March 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979877.

Edgeworth:1899:RSMa

[417] F. Y. Edgeworth. On the representation of statistics by mathematical formulæ. Journal of the Royal Statistical Society, 62(1):125–140, March 1899. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979878.

Bowley:1899:SWUa

[418] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (part i.) agricultural wages. Journal of the Royal Statistical Society, 62(1):140–150, March 1899. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979879.

Parker:1899:NSS

[419] E. H. Parker. A note on some statistics regarding China. Journal of the Royal Statistical Society, 62(1):150– 156, March 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979880.

Coghlan:1899:RCP

[420] T. A. Coghlan. Reply to criticisms on paper, “Deaths in child-birth in New South Wales.”. Journal of the Royal Statistical Society, 62(1):157–159, March 1899. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2979881. See

[396].

Anonymous:1899:CHR

[421] Anonymous. Commercial history and review of 1898. Journal of the Royal Statistical Society, 62(1):159–178, March 1899. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979882.

Sauerbeck:1899:PC

[422] A. Sauerbeck. Prices of commodities in 1898. Journal of the Royal Statistical Society, 62(1):179–193, March 1899. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979883.

Anonymous:1899:FLM

[423] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1898. Journal of the Royal Statistical Society, 62(1):194–198, March 1899. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979884.

Anonymous:1899:EL

[424] Anonymous. English literature in 1898. Journal of the Royal Statistical Society, 62(1):198–200, March 1899. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979885.

Anonymous:1899:NESa

[425] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 62(1):201– 211, March 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979886.

Anonymous:1899:QLAa

[426] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 62(1):212–225, March 1899. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979887.

Anonymous:1899:PR

[427] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 62(1): 226–248, March 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979888.

Yule:1899:ICC

[428] G. Udny Yule. An investigation into the causes of changes in pauperism in England, chiefly during the last two intercensal decades (part i.). Journal of the Royal Statistical Society, 62(2):249–295, June 1899. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979889.

Martineau:1899:SAS

[429] George Martineau. The statistical aspect of the sugar question. Journal of the Royal Statistical Society, 62(2):296–347, June 1899. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979890.

Darwin:1899:SGS

[430] Leonard Darwin. The stability of gold and silver prices in recent years. Journal of the Royal Statistical Society, 62(2): 348–372, June 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979891.

Edgeworth:1899:RSMb

[431] F. Y. Edgeworth. On the representation of statistics by mathematical formulæ. Journal of the Royal Statistical Society, 62(2):373–385, June 1899. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979892.

Anonymous:1899:Ca

[432] Anonymous. The census of 1901. Journal of the Royal Statistical Society, 62 (2):385–389, June 1899. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979893.

Anonymous:1899:AR

[433] Anonymous. The agricultural returns for 1898. Journal of the Royal Statistical Society, 62(2):389–395, June 1899. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979894.

Bowley:1899:SWUb

[434] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (Part III). agricultural wages. Journal of the Royal Statistical Society, 62(2):395–404, June 1899. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979895.

Anonymous:1899:NESb

[435] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 62(2):404–413, June 1899. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979896.

Anonymous:1899:QLAb

[436] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 62(2):413–426, June 1899. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979897.

Anonymous:1899:FYE

[437] Anonymous. For the financial year ended 31st December, 1898, and for the sessional year ending 30th June, 1899, presented at the sixty-fifth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 30th of June, 1899. Journal of the Royal Statistical Society, 62(3):427– 438, September 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979932.

Bateman:1899:PSF

[438] A. E. Bateman. Proceedings of the sixtyfifth Annual General Meeting. Journal of the Royal Statistical Society, 62(3): 439–442, September 1899. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979933.

Hayward:1899:LTTa

[439] T. E. Hayward. On life-tables-their construction and practical application. Journal of the Royal Statistical Society, 62(3):443–488, September 1899. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979934.

Flux:1899:FTS

[440] A. W. Flux. The flag and trade: A summary review of the trade of the Chief colonial empires. Journal of the Royal Statistical Society, 62(3):489– 533, September 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979935.

Anonymous:1899:RSM

[441] Anonymous. On the representation of statistics by mathematical formulæ. Journal of the Royal Statistical Society, 62(3):534–555, September 1899. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979936.

Bowley:1899:SWUc

[442] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (Part IV). agricultural wages. Journal of the Royal Statistical Society, 62(3):555–570, September 1899. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979937.

Anonymous:1899:ARG

[443] Anonymous. Agricultural returns of Great Britain, 1899. Journal of the Royal Statistical Society, 62(3):571–572, September 1899. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979938.

Anonymous:1899:NPI

[444] Anonymous. Notes on the proceedings of the International Statistical Congress in Christiania. Journal of the Royal Statistical Society, 62(3):572–574, September 1899. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979939.

Anonymous:1899:NESc

[445] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 62(3):575–584, September 1899. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979940.

Anonymous:1899:QLAc

[446] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 62(3):584–596, September 1899. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979941.

Anonymous:1899:VIb

[447] Anonymous. Volume information. Journal of the Royal Statistical Society, 62 (4):755–769, December 1899. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979763.

Crawford:1899:NFS

[448] R. F. Crawford. Notes on the food supply of the United Kingdom, Belgium, France, and Germany. Journal of the Royal Statistical Society, 62(4):597– 638, December 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979764.

Wood:1899:SSR

[449] George H. Wood. Some statistics relating to working class progress since 1860. Journal of the Royal Statistical Society, 62(4):639–675, December 1899. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979765.

Anonymous:1899:LF

[450] Anonymous. Lord Farrer. Journal of the Royal Statistical Society, 62(4):676– 677, December 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979766.

Anonymous:1899:SRW

[451] Anonymous. Sir Rawson W. Rawson.

Journal of the Royal Statistical Society, 62(4):677–679, December 1899. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979767.

Anonymous:1899:Cb

[452] Anonymous. The census of 1901. Journal of the Royal Statistical Society, 62 (4):679–683, December 1899. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979768.

Hayward:1899:LTTb

[453] T. E. Hayward. On life-tables: Their construction and practical application. Journal of the Royal Statistical Society, 62(4):683–702, December 1899. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979769.

Everett:1899:GIT

[454] J. D. Everett. On geometrical illustrations of the theory of rent. Journal of the Royal Statistical Society, 62(4):703– 707, December 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979770.

Bowley:1899:SWUd

[455] A. L. Bowley and George Hy. Wood. The statistics of wages in the United Kingdom during the last hundred years. (part v.) printers. Journal of the Royal Statistical Society, 62(4):708–715, December 1899. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979771.

Hooker:1899:BCF

[456] R. H. Hooker. The British and continental Farmer: A comparison. Journal of the Royal Statistical Society, 62(4):715– 723, December 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979772.

Anonymous:1899:ARU

[457] Anonymous. Agricultural returns of the United Kingdom for 1899. Journal of the Royal Statistical Society, 62(4):724– 726, December 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979773.

Anonymous:1899:NESd

[458] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 62(4):726– 738, December 1899. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979774.

Anonymous:1899:QLAd

[459] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 62(4):739–754, December 1899. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979775.